OpenStudy (anonymous):

convert pi/2 radians into degrees

jimthompson5910 (jim_thompson5910):

multiply by 180/pi to convert from radians to degrees

OpenStudy (anonymous):

\[\frac{ \pi }{ 2 } \times \frac{ 180 }{ \pi }\]

OpenStudy (anonymous):

* To convert degrees to radians, multiply by \[\frac{ \pi }{ 180 }\]

OpenStudy (anonymous):

* To convert radians to degrees, multiply by\[ \frac{ 180}{ \pi }\]